A biosensor completely oxidizes the glucose in a blood sample via two-electron redox reaction. What is the concentration (g/L) of glucose in a 0.50 mL blood sample if the sensor requires 0.815 Coulombs?
Answer #1

Charge needed for 1 mole of 2 electron reduction = 2 * 1 F = 2 * 96500 = 1.93 * 105 C

Moles of glucose = Charge used / Charge per mole = 0.815 / (1.93 * 105 ) = 4.223 * 10-6 moles

Mass of glucose = moles * molar mass = 4.223 * 10-6 * 180.16 = 7.608 * 10-4 g

Concentration in g/L = mass in g / Volume in L = 7.608 * 10-4 / (0.5 * 10-3) = 1.522 g/L